0 0
Add a comment Improve this question Transcribed image text
Answer #1

(a)

?S=nCv ln(T2/T1)= 1 * (1.5R) ln(500/298) = 6.454 J/mol K

Sf ( at 500K ) = S (at 298K ) + Delta S = 146.33 J/mol K + 6.454 J/Kmol = 152.784 J/mol K

(b)

?S=nCp ln(T2/T1) = 1* (2.5R) * ln( 500/298) = 10.757 J/mol K

S(at 500K ) = S( at 298 K) + delta S = 10.757 J/mol K + 146.33 J/mol K = 157.087 J/mol K